OpenStudy (anonymous):

Solve: 0 = 4x2 - 8x - 5

OpenStudy (anonymous):

x=5/2,-1/2

OpenStudy (zzr0ck3r):

4x^2-8x-5=0 4x^2+2x-10x-5=0 2x(2x+1)-5(2x+1)=0 (2x+1)(2x-5)=0 2x+1=0 so x=-1/2 2x-5=0 so x = 5/2

OpenStudy (anonymous):

Actually I'm not sure what the equation is... Clarification please. Is it 4x squared or 4 times 2?

OpenStudy (zzr0ck3r):

I think its squared, she has posted a few of these

OpenStudy (zzr0ck3r):

if x was multiply the 8x would not make since

OpenStudy (anonymous):

Ah okay.